OpenStudy (chris215):

Find the average value of f(x) = 1 / (x +1) over the interval [0, 4].

OpenStudy (chris215):

I got 0.40

OpenStudy (mariamikayla):

f(0)= 1/(0+1)= 1/1=1 f(1)=1/(1+1)=1/2= 0.5 f(2)=1/(2+1)=1/3=0.3 f(3)=1/(3+1)=1/4=0.25 f(4)=1/(4+1)=1/5=0.2 Average is ( 1+0.5+0.3+0.25+0.2)/5= 2/25/4= 0.45

OpenStudy (chris215):

ohh hm the closest answer is 0.40 so its probably correct thank you!

OpenStudy (mariamikayla):

you're welcome :D

OpenStudy (robtobey2):

http://tutorial.math.lamar.edu/Classes/CalcI/AvgFcnValue.aspx \[\frac{\int\limits_0^4 \frac{1}{x+1} \, dx}{4+0}=\frac{\log (5)}{4}=0.402359\text{...} \]